OpenStudy (anonymous):

what's the derivative of (sin^(5)a)(cos^(5)a)?

OpenStudy (anonymous):

is a constant or variable??

OpenStudy (anonymous):

a is a variable

OpenStudy (anonymous):

if constant then zero for a is variable 5*sin^4a*cos^5a*cosa - 5*cos^4a*sin^5a*sina

OpenStudy (anonymous):

\[(\sin ^{5}a)(\cos ^{5}a)\]

OpenStudy (anonymous):

\[5(\sin^4x)(\cos^6x)-5(\cos^4x)(\sin^6x)\]

OpenStudy (anonymous):

Substitute a for x

OpenStudy (anonymous):

Use the product rule.